OpenStudy (anonymous):

I've been at this problem for like 20 minutes. help?! 8 – (3 + b) = b – 9 a.7 b.8 c.no solution d.identity

OpenStudy (lgbasallote):

8 - 3 - b = b - 9 8 - 3 +9 = 2b 14 = 2b b= 7

OpenStudy (anonymous):

oh, I see what I did wrong now. lol thanks!